# Entropic curvature not comparable to other curvatures – or is it?

### v1

@@ 1 Introduction
In [8], Erbar and Maas introduced in 2012 entropic curvature for finite Markov chains $(X,Q,\pi)$ with transition probabilities $Q:X\times X\to[0,1]$ and steady state $\pi:X\to[0,1]$ by way of a particular mean $\theta_{log}$, called the logarithmic mean (see also [23] for related work). In this paper, we will always assume that our finite Markov chains are irreducible and reversible, and the transition probabilities $Q$ induce a canonical weighted graph structure on $X$, where edges correspond to positive transition probabilities. Conversely, any weighted graph $G=(V,E)$ with vertex measure $m:V\to(0,\infty)$ and symmetric edge weights $w:V\times V\to[0,\infty)$ give rise to a canonical finite irreducible, reversible Markov chain with $Q(x,y)=\frac{w_{xy}}{m(x)}$ for $x\sim y$. For details about this, see Subsection 2.1.
The notion of entropic curvature is non-local in nature and generalizes naturally to other means $\theta:[0,\infty)\times[0,\times)\to[0,\infty)$, like the arithmetic or the geometric mean. In the case of the arithmetic mean, the resulting curvature is closely related to Bakry-Émery curvature. This relation was well-known and mentioned at various places, and was further manifested by a modified $\Gamma$ calculus (given in [15]) for entropic curvature and more general $\theta$-curvatures. The $\Gamma$ operators are defined locally and are crucial in the definition of the Bakry-Émery curvature, which is defined locally on the vertices of a finite Markov chain and which includes an additional dimension parameter. This approach via modified $\Gamma$ operators motivates also the introduction of a dimension parameter into entropic curvature and into more general $\theta$-curvatures. The parameters $K\in\mathbb{R}$ and $N\in(0,\infty]$ in the $CD_{\theta}(K,N)$-property of a finite Markov chain $(X,Q,\pi)$ represent lower and upper bounds on global curvature and dimension, respectively. Since we will make also use of the $\Gamma$ formalism in this paper, we provide a brief overview about it in Subsection 2.4.
@@ 1.1 Our results
In fact, the third pair $(K,N)$ in the theorem is deduced by the first two pairs by using the trade-off argument given in [15, Remark 3.11] (with $\alpha=1/2$). The first two pairs are the results presented in Theorems 3.4 and 3.5 in Subsections 3.2 and 3.3, respectively. Note that $CD_{ent}(0,|S|-0.07|S\backslash S_{2}|)$ has a slightly better dimension parameter than the corresponding Bakry-Émery curvature-dimension condition $CD_{a}(0,|S|)$ for these abelian Cayley graph Markov chains (see Remark 3.6).
Non-negativity of entropic curvature for these abelian Cayley graph Markov chains was already proved by [8, Proposition 5.4]. In fact, they formulated this result for Markov chains with a mapping representation, which is equivalent to the abelian Cayley graph structure (see Theorem 3.3 in Subsection 3.1). Moreover, these particular Markov chains have also non-negative Bakry-Émery and Ollivier Ricci curvatures (see, e.g., [5] and references therein). The above theorem stating strictly positive entropic curvature for these Markov chains is surprising, since cycles of large enough diameter have always vanishing Bakry-Émery and Ollivier Ricci curvature. Moreover, we cannot deduce vanishing first homology from strictly positive entropic curvature (in contrast to Bakry-Émery and Forman curvature, see [16, 27, 13]).
@@
It was already shown in [23, Theorem 4.1] that the entropic curvature of every finite Markov chain is greater than $\infty$- ∞. Our theorem provides an explicit lower bound and also holds under the stronger finite dimension assumption. The main ideas of our theorem were subsequently used to derive a lower bound for the intertwining curvature in [28].
@@
[…]
@@ 1.2 Previous work on discrete curvature
for suitable norms and gradients. The reason that the discrete Ricci curvature notions are different (and not even comparable), is the difference in measuring the gradient, and the lack of a chain rule for the gradient and Laplacian. An overview on discrete Ricci curvature is given in the book [30]. For work on Bakry Emery Ricci curvature, see e.g. [11, 40, 20, 17]. For work on Ollivier Ricci curvature, see e.g. [35, 33, 19, 12]. In this article, we are particularly interested in entropic Ricci curvature. For previous work, see [10, 9, 6, 7, 23, 15]. In contrast to the entropic curvature by Erbar, Maas and Mielke, there is also a different notion of entropic curvature introduced by Samson and based on Schrödinger bridges for the $\ell_{1}$-Wasserstein distance, see [39, 37, 36]. Recently, Ricci curvature has been proven to be a useful tool in machine learning, and in particular, in graph neural networks [43, 31, 44].
@@ 2.1 Markov chains and means
For curvature notions, in particular entropic curvature, we need the notion of a mean, as introduced in [22, 8]. The relevant mean for entropic curvature is called logarithmic mean.
@@ 2.2 Relevant properties of the logarithmic mean
The function $b$ is a crucial ingredient for our analysis later in Section 3. The function $b_{0}$ will be relevant in Section 4. In addition to its nonnegativity, which can be seen from [8, Lemma 2.2], we present useful properties and estimates for the function $b$ in the following proposition. (Note that in the case of the arithmetic mean $\theta=\theta_{a}$, the corresponding function $b$, defined in (2) would vanish identically.)
@@ 2.3 Entropic curvature
In analogy to the Ricci curvature notion by Sturm [41, 42] and Lott-Villani [21] for metric measure spaces, entropic curvature is defined via a convexity property along geodesics with respect to suitably modified Wasserstein metric, adapted to the discrete setting of a finite Markov chain. For more details, we refer readers to the original papers [22, 8]. In this paper, we focus on an equivalent infinitesimal curvature formulation of entropic curvature (see Definition 2.5 below), which is an integrated Bochner-type inequality and involves two operators, denoted by $\mathcal{A}_{\rho}$ and $\mathcal{B}_{\rho}$ for all $\rho\in\mathcal{P}_{*}(X)$, where
@@
Moreover, we have for any mean (see [8, Lemma 2.2]):
@@
We use the following infinitesimal definition of entropic curvature (combining [8, Definition 1.1] and [8, Theorem 4.5]):
@@ 2.4 Adapted $\Gamma$ calculus and $\theta$-curvatures
Let $(X,Q,\pi)$ be a finite Markov chain and $\theta:[0,\infty)\times[0,\infty)\to[0,\infty)$ be a mean. In analogy to the Bakry-Émery $\Gamma$ calculus (see [1]), we introduce for every $\rho\in(0,\infty)^{X}$ the following $\rho$-Laplacian and modified bilinear operators $\Gamma_{\rho},\Gamma_{2,\rho}$ (see [15, Definition 3.1]):
@@
Note that the definition of $\Gamma_{\rho}$ involves the $\rho$-Laplacian, while $\Gamma_{2,\rho}$ involves the standard Laplacian. We like to emphasize that the operators $\Gamma_{\rho}$ and $\Gamma_{2,\rho}$ reduce to the classical $\Gamma$ and $\Gamma_{2}$ operators, involving only the standard Laplacian, if we choose $\theta=\theta_{a}$, that is, the arithmetic mean, since in this case we have $12\partial_{1}\theta(r,s)=$ 1 for all $r,s\geq 0$. It was shown in [15, Section 3.1] that we have
@@
We can think of (13) as a (non-local) integrated curvature-dimension inequality against the probability density $\rho$, and we recover Definition 2.5 via the special choices $\theta=\theta_{log}$ and $N=\infty$. Moreover, as stated in [15, Lemma 4.1], the best lower $\theta_{a}$-curvature bound for $(X,Q,\pi)$ is given by $\min_{x\in X}K_{N}(x)$, where $K_{N}(x)$ is the (local) Bakry-Émery curvature at $x\in X$ introduced in Definition 2.6.
@@ 3.1 Special mapping representations and abelian Cayley graphs
In preparation for the following subsections, we recall certain facts from [8] for Markov chains $(X,Q,\pi)$ with mapping representations and provide some further notation. We assume the mapping representations to satisfy the extra conditions (i) and (ii) mentioned in Theorem 3.3. Therefore, we have $c(x,\delta)=c(\delta)$ and $c(\delta)=c(\delta^{-1})$, and that $\pi$ is constant, so we can set $\pi\equiv\pi_{0}$. We have with the notation given in [8, pages 1024-1026],
@@ 3.3 Lower curvature bound for finite dimension
The proof of Theorem 3.5 can be adapted to Bakry-Émery curvature with the only difference being $N=|S|$ (that is, using $b\equiv 0$ in the case $\theta=\theta_{a}$ in the proof). This means an abelian Cayley graph with generating set $S$ always satisfies the Bakry-Émery curvature condition $CD_{a}(0,|S|)$. In the special case that all reduced relations of the generators except for the commutator relations have length $\geq 5$, this fact is also confirmed by Theorem 9.1 in [4]. Furthermore, abelian Cayley graphs are always Ricci flat graphs but not vice versa (see [18] for more details). Ricci flat graphs were introduced by [3] and they are $d$-regular graphs satisfying also $CD_{a}(0,d)$, as mentioned in [24]. Moreover, it was shown in [2] that Ricci flat graphs satisfy $CDE(0,d)$.
@@
Let $C_{n}$ be the cycle of length $n$, equipped with the simple random walk $Q(x,y)=\frac{1}{2}$ for adjacent vertices $x,y$. In this subsection, we improve the trivial upper bound $K_{ent}\leq\lambda_{1}=2\sin^{2}\left(\frac{\pi}{n}\right)\leq\frac{2\pi^{2}}{n^{2}}$, derived from the Lichnerowicz inequality (see, e.g., [15, Theorem 1.3]) to $K\lesssim n^{-4}\log(n)^{2}$.

### v3

@@ 1 Introduction
In [EM-12], Erbar and Maas introduced in 2012 entropic curvature for finite Markov chains $(X,Q,\pi)$ with transition probabilities $Q:X\times X\to[0,1]$ and steady state $\pi:X\to[0,1]$ by way of a particular mean $\theta_{log}$, called the logarithmic mean (see also [Mi-13] for related work). In this paper, we will always assume that our finite Markov chains are irreducible and reversible, and the transition probabilities $Q$ induce a canonical weighted graph structure on $X$, where edges correspond to positive transition probabilities. Conversely, any weighted graph $G=(V,E)$ with vertex measure $m:V\to(0,\infty)$ and symmetric edge weights $w:V\times V\to[0,\infty)$ give rise to a canonical finite irreducible, reversible Markov chain with $Q(x,y)=\frac{w_{xy}}{m(x)}$ for $x\sim y$. For details about this, see Subsection 2.1.
The notion of entropic curvature is non-local in nature and generalizes naturally to other means $\theta:[0,\infty)\times[0,\times)\to[0,\infty)$, like the arithmetic or the geometric mean. In the case of the arithmetic mean, the resulting curvature is closely related to Bakry-Émery curvature. This relation was well-known and mentioned at various places, and was further manifested by a modified $\Gamma$ calculus (given in [KLMP-23]) for entropic curvature and more general $\theta$-curvatures. The $\Gamma$ operators are defined locally and are crucial in the definition of the Bakry-Émery curvature, which is defined locally on the vertices of a finite Markov chain and which includes an additional dimension parameter. This approach via modified $\Gamma$ operators motivates also the introduction of a dimension parameter into entropic curvature and into more general $\theta$-curvatures. The parameters $K\in\mathbb{R}$ and $N\in(0,\infty]$ in the $CD_{\theta}(K,N)$-property of a finite Markov chain $(X,Q,\pi)$ represent lower and upper bounds on global curvature and dimension, respectively. Since we will make also use of the $\Gamma$ formalism in this paper, we provide a brief overview about it in Subsection 2.4.
@@ 1.1 Our results
In fact, the third pair $(K,N)$ in the theorem is deduced by the first two pairs by using the trade-off argument given in [KLMP-23, Remark 3.11] (with $\alpha=1/2$). The first two pairs are the results presented in Theorems 3.4 and 3.5 in Subsections 3.2 and 3.3, respectively. Note that $CD_{ent}(0,|S|-0.07|S\backslash S_{2}|)$ has a slightly better dimension parameter than the corresponding Bakry-Émery curvature-dimension condition $CD_{a}(0,|S|)$ for these abelian Cayley graph Markov chains (see Remark 3.6).
Non-negativity of entropic curvature for these abelian Cayley graph Markov chains was already proved by [EM-12, Proposition 5.4]. In fact, they formulated this result for Markov chains with certain mapping representations, which is equivalent to the abelian Cayley graph structure (see Theorem 3.3 in Subsection 3.1). The concept of a mapping representation, as introduced in [EM-12, Definition 5.2], follows in spirit the conditions given in the earlier paper [CDPP-09]. This earlier paper studies a Bochner-Bakry-Émery approach to the Modified Logarithmic Sobolev Inequality (MLSI) and considers particular examples: various birth-death processes and zero range processes, and Bernoulli-Laplace models which are equivalent to random walks on Johnson graphs.
Moreover, abelian Cayley graph Markov chains have also non-negative Bakry-Émery and Ollivier Ricci curvatures (see, e.g., [cushing2021curvatures] and references therein). The above theorem stating strictly positive entropic curvature for these Markov chains is surprising, since cycles of large enough diameter have always vanishing Bakry-Émery and Ollivier Ricci curvature. Moreover, we cannot deduce vanishing first homology from strictly positive entropic curvature (in contrast to Bakry-Émery and Forman curvature, see [KMY-21, MR-20, For-03]).
@@
It was already shown in [Mi-13, Theorem 4.1] that the entropic curvature of every finite Markov chain is greater than $\infty$- ∞. Our theorem provides an explicit lower bound and also holds under the stronger finite dimension assumption. The main ideas of our theorem were subsequently used to derive a lower bound for the intertwining curvature in [MWZ-24].
@@
$\sup_{x\sim y}\left|\log\frac{\widetilde{w}_{xy}}{w_{xy}}\right|\leq\epsilon\quad\text{and}\quad\sup_{x\in X}\left|\log\frac{\widetilde{m}(x)}{m(x)}\right|\leq\epsilon.$
@@ 1.2 Previous work on discrete curvature
for suitable norms and gradients. The reason that the discrete Ricci curvature notions are different (and not even comparable), is the difference in measuring the gradient, and the lack of a chain rule for the gradient and Laplacian. An overview on discrete Ricci curvature is given in the book [najman2017modern]. For work on Bakry Emery Ricci curvature, see e.g. [fathi2015curvature, schmuckenschlager1998curvature, lin2010ricci, KKRT-16]. For work on Ollivier Ricci curvature, see e.g. [ollivier2007ricci, ollivier2009ricci, lin2011ricci, fathi2022discrete]. In this article, we are particularly interested in entropic Ricci curvature. For previous work, see [FM-16, erbar2017ricci, erbar2018poincare, Erbar2019EntropicCA, Mi-13, KLMP-23]. In contrast to the entropic curvature by Erbar, Maas and Mielke, there is also a different notion of entropic curvature introduced by Samson and based on Schrödinger bridges for the $\ell_{1}$-Wasserstein distance, see [Sam-22, rapaport2023criteria, pedrotti2023contractive]. Recently, Ricci curvature has been proven to be a useful tool in machine learning, and in particular, in graph neural networks [topping2021understanding, nguyen2023revisiting, ye2019curvature].
@@ 2.1 Markov chains and means
For curvature notions, in particular entropic curvature, we need the notion of a mean, as introduced in [Ma-11, EM-12]. The relevant mean for entropic curvature is called logarithmic mean.
@@ 2.2 Relevant properties of the logarithmic mean
The function $b$ is a crucial ingredient for our analysis later in Section 3. The function $b_{0}$ will be relevant in Section 4. In addition to its nonnegativity, which can be seen from [EM-12, Lemma 2.2], we present useful properties and estimates for the function $b$ in the following proposition. (Note that in the case of the arithmetic mean $\theta=\theta_{a}$, the corresponding function $b$, defined in (2) would vanish identically.)
@@ 2.3 Entropic curvature
In analogy to the Ricci curvature notion by Sturm [St-06-i, St-06-ii] and Lott-Villani [LV-09] for metric measure spaces, entropic curvature is defined via a convexity property along geodesics with respect to suitably modified Wasserstein metric, adapted to the discrete setting of a finite Markov chain. For more details, we refer readers to the original papers [Ma-11, EM-12]. In this paper, we focus on an equivalent infinitesimal curvature formulation of entropic curvature (see Definition 2.5 below), which is an integrated Bochner-type inequality and involves two operators, denoted by $\mathcal{A}_{\rho}$ and $\mathcal{B}_{\rho}$ for all $\rho\in\mathcal{P}_{*}(X)$, where
@@
Moreover, we have for any mean (see [EM-12, Lemma 2.2]):
@@
We use the following infinitesimal definition of entropic curvature (combining [EM-12, Definition 1.1] and [EM-12, Theorem 4.5]):
@@ 2.4 Adapted $\Gamma$ calculus and $\theta$-curvatures
Let $(X,Q,\pi)$ be a finite Markov chain and $\theta:[0,\infty)\times[0,\infty)\to[0,\infty)$ be a mean. In analogy to the Bakry-Émery $\Gamma$ calculus (see [BE-85]), we introduce for every $\rho\in(0,\infty)^{X}$ the following $\rho$-Laplacian and modified bilinear operators $\Gamma_{\rho},\Gamma_{2,\rho}$ (see [KLMP-23, Definition 3.1]):
@@
Note that the definition of $\Gamma_{\rho}$ involves the $\rho$-Laplacian, while $\Gamma_{2,\rho}$ involves the standard Laplacian. We like to emphasize that the operators $\Gamma_{\rho}$ and $\Gamma_{2,\rho}$ reduce to the classical $\Gamma$ and $\Gamma_{2}$ operators, involving only the standard Laplacian, if we choose $\theta=\theta_{a}$, that is, the arithmetic mean, since in this case we have $12\partial_{1}\theta(r,s)=$ 1 for all $r,s\geq 0$. It was shown in [KLMP-23, Section 3.1] that we have
@@
We can think of (13) as a (non-local) integrated curvature-dimension inequality against the probability density $\rho$, and we recover Definition 2.5 via the special choices $\theta=\theta_{log}$ and $N=\infty$. Moreover, as stated in [KLMP-23, Lemma 4.1], the best lower $\theta_{a}$-curvature bound for $(X,Q,\pi)$ is given by $\min_{x\in X}K_{N}(x)$, where $K_{N}(x)$ is the (local) Bakry-Émery curvature at $x\in X$ introduced in Definition 2.6.
@@ 3.1 Special mapping representations and abelian Cayley graphs
In preparation for the following subsections, we recall certain facts from [EM-12] for Markov chains $(X,Q,\pi)$ with mapping representations and provide some further notation. We assume the mapping representations to satisfy the extra conditions (i) and (ii) mentioned in Theorem 3.3. Therefore, we have $c(x,\delta)=c(\delta)$ and $c(\delta)=c(\delta^{-1})$, and that $\pi$ is constant, so we can set $\pi\equiv\pi_{0}$. We have with the notation given in [EM-12, pages 1024-1026],
@@ 3.3 Lower curvature bound for finite dimension
The proof of Theorem 3.5 can be adapted to Bakry-Émery curvature with the only difference being $N=|S|$ (that is, using $b\equiv 0$ in the case $\theta=\theta_{a}$ in the proof). This means an abelian Cayley graph with generating set $S$ always satisfies the Bakry-Émery curvature condition $CD_{a}(0,|S|)$. In the special case that all reduced relations of the generators except for the commutator relations have length $\geq 5$, this fact is also confirmed by Theorem 9.1 in [CLP-19]. Furthermore, abelian Cayley graphs are always Ricci flat graphs but not vice versa (see [LL-24] for more details). Ricci flat graphs were introduced by [CY-96] and they are $d$-regular graphs satisfying also $CD_{a}(0,d)$, as mentioned in [munch2018li]. Moreover, it was shown in [bauer2015li] that Ricci flat graphs satisfy $CDE(0,d)$.
@@
Let $C_{n}$ be the cycle of length $n$, equipped with the simple random walk $Q(x,y)=\frac{1}{2}$ for adjacent vertices $x,y$. In this subsection, we improve the trivial upper bound $K_{ent}\leq\lambda_{1}=2\sin^{2}\left(\frac{\pi}{n}\right)\leq\frac{2\pi^{2}}{n^{2}}$, derived from the Lichnerowicz inequality (see, e.g., [KLMP-23, Theorem 1.3]) to $K\lesssim n^{-4}\log(n)^{2}$.
@@
3.5 Abelian Cayley graphs and Ricci flatness
Abelian Cayley graphs belong to the class of Ricci flat graphs, which were introduced in [CY-96]. The following definition presents special versions of this notion:
Definition 3.9 (see [cushing2021curvatures, Def. 3.1]).
Let $G=(V,E)$ be a $d$-regular graph. A vertex $x\in V$ is Ricci flat if there exist maps $\eta_{j}:B_{1}(x)\to V$, $1\leq j\leq d$, with the following properties:
  • (i)
    $\eta_{j}(y)\sim y$ for all $y\in B_{1}(x)$,
  • (ii)
    $\eta_{i}(y)\neq\eta_{j}(y)$ for all $y\in B_{1}(x)$ and $i\neq j$,
  • (iii)
    $\bigcup_{j}\eta_{j}(\eta_{i}(x))=\bigcup_{j}\eta_{i}(\eta_{j}(x))$ for all $i$.
The additional properties
$(R)$ $\quad\eta_{i}^{2}(x)$ $=x\quad\text{for all $i$,}$
$(S)$ $\quad\eta_{j}(\eta_{i}(x))$ $=\eta_{i}(\eta_{j}(x))\quad\text{for all $i,j$,}$
are called reflexivity and symmetry, respectively.
If there exists a family of maps $\eta_{j}$ for a given vertex $x\in V$ satisfying (R) or (S) in addition to (i)–(iii), we say that $x$ is $(R)$-Ricci flat or $(S)$-Ricci flat, respectively. If there exists a family of maps $\eta_{j}$ satisfying (i)–(iii) and (R) and (S) simultaneously, we say that $x$ is $(RS)$-Ricci flat.
The (S)-Ricci flatness definition is similar in spirit to the special mapping representation considered in Theorem 3.2 with the main difference that Ricci flatness is a local property. It is known that Ricci flatness implies both non-negative Ollivier Ricci curvature and non-negative Bakry-Émery curvature (see, e.g., [cushing2021curvatures] and the references therein). It is not difficult to see that Ricci flatness also implies non-negative Ollivier sectional curvature (see Definition 5.1 below).
It would be interesting to investigate whether Ricci flatness also implies non-negative entropic curvature. Let us discuss some potential naive strategies to tackle this question.
Note that all abelian Cayley graphs are (S)-Ricci flat, but not vice versa. Examples of such graphs can be found in [LL-24]. Optimistic readers may hope that every Ricci flat graph has an abelian Cayley graph as a covering. A sujective graph homomorphism $\pi:\tilde{G}\to G$ is called a covering if it is locally bijective, that is one-balls $B_{1}(x)$ of $\tilde{G}$ are bijectively mapped onto one-balls $B_{1}(\pi(x))$ of $G$. A counterexample for the above hope is the graph in Figure 3 (the complement of $C_{4}\dot{\cup}C_{5}$) together with Lemma 3.10. Since this graph is highly connected, we suspect that it has positive entropic curvature.
Figure 3: A Ricci flat graph which is not (S)-Ricci flat, and therefore not an abelian Cayley graph (vertices labelled “F” and “R” are only Ricci-flat and only (R)-Ricci flat, respectively). The picture was generated via the Graph Curvature Calculator [CKLLS-22].
Lemma 3.10.
Let $\pi:\tilde{G}\to G$ be a covering. Let $x$ be a vertex in $G$ and $\tilde{x}\in\pi^{-1}(\{x\})\in\tilde{G}$. If $\tilde{x}$ is (S)-Ricci flat in $\tilde{G}$, then $x$ is also (S)-Ricci flat in $G$.
Proof.
Let $\tilde{G}=(\tilde{V},\tilde{E})$ and $G=(V,E)$. Assume that $\tilde{x}\in\tilde{V}$ is $(S)$-Ricci flat. Then $\pi|_{B_{1}(\tilde{x})}:B_{1}(\tilde{x})\to B_{1}(x)$ is bijective with inverse $\rho:B_{1}(x)\to B_{1}(\tilde{x})$. Let $\tilde{\eta}_{j}:B_{1}(\tilde{x})\to\tilde{V}$ be as in Definition 3.9 and
$\eta_{j}=\pi\circ\tilde{\eta}_{j}\circ\rho:B_{1}(x)\to V.$
Since both $\rho$ and $\pi$ are graph homomorphisms and $\tilde{\eta}_{j}(\tilde{y})\sim\tilde{y}$ for all $\tilde{y}\in B_{1}(\tilde{x})$, we have $\eta_{j}(y)\sim y$ for all $y\in B_{1}(x)$. This shows (i).
Property (ii) follows from the local bijectivity of $\pi$ around $\rho(y)$.
Property (S) (and therefore also (iii)) follows via the following identities:
$\eta_{j}(\eta_{i}(x))$ $=\pi\circ\tilde{\eta}_{j}\circ(\rho\circ\pi)\circ\tilde{\eta}_{i}\circ\rho(x)$
$=\pi\circ\tilde{\eta}_{j}\circ\tilde{\eta}_{i}\circ\rho(x)$
$=\pi\circ\tilde{\eta}_{i}\circ\tilde{\eta}_{j}\circ\rho(x)$
$=\eta_{i}(\eta_{j}(x)).$